OpenStudy (magentama202):

QUICK ANSWER NEEDED PLZ. I don't know how to create a table from these equations... on part B, would also like help with C: Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 4−x and y = 2x + 3 intersect are the solutions of the equation 4−x = 2x + 3. Part B: Make tables to find the solution to 4−x = 2x + 3. Take the integer values of x between −3 and 3. Part C: How can you solve the equation 4−x = 2x + 3 graphically?

OpenStudy (magentama202):

@mjdennis can you help me?

OpenStudy (mjdennis):

x 4-x= 2x+3= =============== -3 4--3 2(-3)+3 -2 ______ ________ etc. That is what they are aiming for, in part B, except you finish the table and finish the math. Start on that while I look at part C

OpenStudy (magentama202):

Ok one sec I just got back.. thank you soo much though!!!

OpenStudy (magentama202):

what does 4--3 mean???

OpenStudy (magentama202):

Oh, you were inserting the (-3) into the equation got it.

OpenStudy (mjdennis):

Part C, take the table from part B. It is built of two sets of ordered pairs (x,4−x) is the first and second columns. Or, we can call them (x,y) because we were given y=4−x. Can you plot (x,y) values?

OpenStudy (mjdennis):

Yep, moved too fast, sorry: 4-(-3)

OpenStudy (magentama202):

x 4-x= 2x+3= =============== -3 7 -3 4-(x) 2(x)+3 -2 6 -1 -1 5 1 Sorry this took so long not my strong subject...

OpenStudy (magentama202):

is this correct??

OpenStudy (mjdennis):

Good so far, keep going. Remeber, the question says to try "integer values of x between −3 and 3"